triangle ABC is similar to triangle XYZ. calculate the length of side YZ.
Triangle ABC
BA-4
AC-4
BC-6
Triangle XYZ
YX-6
XZ-6
YZ-?
a.6
b.8
c.9
d.12

Answer by richwmiller(17219)

4/6=6/x
2/3=6/x
2x=18
x=9